Mr. Shen wished to enter the Orange Kite Competition with an innovative kite of his own design. He did tests that showed an irregular quadrilateral design was best.​ The longest side​ (A) was twice the length of the shortest side​ (C). Side B was one and​ one-half times the shortest side​ (C). Side D needed to be 3
​" left 3 inches right longer than side C. If the competition required that the perimeter of your kite must be 69​",

what were the dimensions of Mr.​ Shen's kite?

  An irregular quadrilateral with sides labeled A, B, C, and D.
Length of side A equals
​"

Answer by Theo(13342)

A is the longest side
C is the shortest side
A = 2C
B = 1.5C]
C = C
D = C + 3 if i understood you correctly.

A + B + C + D = 69
convert all to C to get:
2C + 1.5C + C + C + 3 = 69
combine like terms to get:
5.5C + 3 = 69
subtract 3 from both sides to get:
5.5C = 66
solve for C to get:
C = 66/5.5 = 12

you get:
A = 24
B = 18
C = 12
D = 15

sum of all sides is 24 + 18 + 12 + 15 = 69
this confirms the solution is correct.
solution is side A = 24, B = 18, C = 12, D = 15.